OpenStudy (anonymous):

can somone please help!!! 90(x - 2) = C???

OpenStudy (anonymous):

@JamesJ

OpenStudy (anonymous):

@anas2000

OpenStudy (anonymous):

@SmokeysTheName

OpenStudy (texaschic101):

90(x - 2) = C 90x - 180 = C 90x = C + 180 x = (C + 180)/90 I guess that is how to do it ???

OpenStudy (whpalmer4):

Always good to check one's answers: \[90(x-2) = C\]\[90(\frac{C+180}{90} - 2) = C\]\[\frac{90C}{90} + \frac{90*180}{90} - 90*2 = C\]\[C + 180 - 180 = C\]\[C=C\checkmark\]